The concentration of acetic acid (HC2H3O2) is 0.1 M and has a density of 0.9987 g/mL. Calculate the mass percent.
Answer #1

0.1 moles of acetic acid is present in 1 L of the solution

Mass of the solution= density* volume of the solution

Mass of the solution= 0.9987 g/ml *1000 ml

Mass of the solution= 998.7 g

Mass of acetic acid ( solute) = 0.1 mole* 60 g/mol = 6 g

Mass of the solvent= 998.7 -6 =992.7 g

Mass percent= mass of the solute*100/ mass of the solution

Mass percent= 6*100/998.7 = 0.6